three times the sum of a number and 4 is the same as 18 more than the number
Answer by mananth(16946)

three times the sum of a number and 4 is the same as 18 more than the number
..
let the number be x
3x+4
x+18
...
3x+4 = x+18
3x-x=18-4
2x=14
/2
x= 7
...
m.ananth@hotmail.ca